# Algebraic K-theory of finite algebras over higher local fields

### v2

@@ 1.5. Notations and conventions
We use the language of $\infty$-categories developed in [Lur09], and the theory of higher algebra developed in [Lur17], for convenience. We denote the category of spectra by $\operatorname{Sp}$, and category of spaces by $\mathscr{S}$. And for an $\mathbb{E}_{n}$-monoidal category $\mathcal{C}$, $n\geq 2$, the category of $\mathbb{E}_{m}$-algebras over $\mathcal{C}$ is denoted by $\operatorname{Alg}_{\mathbb{E}_{m}}(\mathcal{C})$, where $0\leq m\leq n$. For $R\in\operatorname{Alg}_{\mathbb{E}_{n}}(\operatorname{Sp})$, we use $\operatorname{Alg}^{(m)}_{R}$ to denote $\operatorname{Alg}_{\mathbb{E}_{m}}(\operatorname{Mod}_{R})$. When $m=1,$ we write $\operatorname{Alg}_{R}^{(1)}$ as $\operatorname{Alg}_{R}$, when $m=\infty$, we write $\operatorname{Alg}_{R}^{(\infty)}$ as $\operatorname{CAlg}_{R}$. We use the symbol $\otimes^{\mathbb{E}_{m}}_{R}$ to indicate the tensor product in $\operatorname{Alg}^{(m)}_{R}$. When $R$ is the unit $\mathbb{S}\in\operatorname{Sp}$, we simply omit $R$. We refer to an object in $\operatorname{Alg}^{(m)}_{R}$ as an $\mathbb{E}_{m}$-$R$-algebra. When $m=1$, we simply refer to an object in $\operatorname{Alg}_{R}$ as an $R$-algebra. When $m=\infty$, we refer to an object in $\operatorname{CAlg}_{R}$ as a commutative $R$-algebra. When $R=\mathbb{S}$, we simply omit the word ’$R$’, or we call it $\mathbb{E}_{m}$-ring spectrum. A (discrete) ring simply indicate a classical ring. Let $R$ be a ring, we use $\otimes^{\mathbb{L}}$ to denote the derived tensor product. We use $\Lambda_{R}(x_{1},\ldots,x_{n})$ to denote the exterior algebra over $R$ generated by $x_{1},\ldots,x_{n}$. And we use $\Gamma_{R}(x_{1},\ldots,x_{n})$ to denote the divided power algebra generated by $x_{1},\ldots,x_{n}$, which sometimes we also write as $R\langle x_{1},\ldots,x_{n}\rangle$. For a stable $\infty$-category $\mathcal{C}$ with a $t$-structure, we use $\tau_{\leq n},\tau_{\geq n}$ to denote the truncation functors. When $\mathcal{C}=\operatorname{Sp}$, $X\in\operatorname{Sp}_{\leq n}$, if the localization $X\to\tau_{\leq n}X$ is an equivalence, and $X\in\operatorname{Sp}_{\geq m}$, if $\tau_{\geq m}X\to X$ is an equivalence. When $X\in\operatorname{Sp}_{\leq n}$, for some $n$, we say $X$ is bounded above, and $X\in\operatorname{Sp}_{\geq m}$, for some $m$, we say that $X$ is bounded below. If there are $a,b$ such that $X\in\operatorname{Sp}_{[a,b]}=\operatorname{Sp}_{\geq a}\cap\operatorname{Sp}_{\leq b}$, then $X$ is said to be bounded. Let $\mathcal{C}$ be a symmetrical monoidal category, and $A\in\operatorname{Alg}(\mathcal{C})$. We use [NS18, Definition III.2.3], to define $\operatorname{THH}(A)$, for instance, $A\in\operatorname{Alg}$, $\operatorname{THH}(A)=A\otimes_{A\otimes A^{\textup{op}}}A$.
@@ 1.6. Acknowledgments
The author would like to thank his advisor Guozhen Wang for carefully reading the early drafts. The author would like to thank Chris Brav for helping to improve the writing of this article. The author would like especially to thank Lars Hesselholt, for suggesting the author Proposition 4.9, and telling the author Proposition 4.19. The author would like to thank Oliver Braunling and John Rognes for comments. Finally, the author thanks Jingbang Guo and Wei Yang, for useful conversations.
@@ 2. Truncated polynomials and infinitesimal thickening
We define $\mathbb{S}[x]:=\Sigma^{\infty}_{+}\mathbb{N}.$ In a same fashion, we may consider the truncated version. For $e\in\mathbb{N}_{\geq 1}$, $M_{e}:=\{1,x,\ldots,x^{e-1}\}\simeq\mathbb{N}/[e,\infty)$ where the multiplication is defined in an obvious way. The multiplication of $M_{e},\mathbb{N}$ are strictly commutative. Unwinding the definitions in [Lur17, Definition 2.4.2.1, Remark 2.4.2.2], we learn that $\mathbb{N},M_{e}$ are commutative monoid objects in $\mathscr{S}$, and hence commutative algebra objects in $\mathscr{S}$. Let $\mathbb{S}[x]/x^{e}:=\Sigma^{\infty}_{+}M_{e}\in\operatorname{Sp}.$ There are $\mathbb{E}_{\infty}$-ring maps
@@
$\mathbb{N}\to\mathbb{N}/[e,\infty)$
@@
$\mathbb{N}/[e,\infty)\to\mathbb{N}/[1,\infty)$
@@
The underlying spectrum of $\mathbb{S}[x]$ is nothing but a direct sum of $\mathbb{S},$ and this is also true for $\mathbb{S}[x]/x^{e},$ using the multiplication on $\mathbb{N},M_{e},$ we can get
@@
In general, let $M$ be a commutative monoid, one can define $A[M]:=A\otimes\Sigma^{\infty}_{+}M$, for any spectrum $A.$ For example, $M$ is a sub monoid of ×$\mathbb{N}^{\times m}$, or a quotient monoid of ×$\mathbb{N}^{\times m}.$ In a upcoming paper, we will consider $\operatorname{BP}\langle n\rangle[x,y]/(xy)$.
@@
There is an $\mathbb{E}_{3}$-algebra $A$, and $A$ is the image of an $\mathbb{E}_{3}$-$\operatorname{MU}[x]$-algebra form of $\operatorname{BP}\langle n\rangle[x]/x^{e}$ under the forgetful functor.
@@
We denote $A$ to be the $\mathbb{E}_{3}$-algebra in Proposition 2.7, $B$ to be the underlying $\mathbb{E}_{3}$-algebra of the $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra $\operatorname{BP}\langle n\rangle$ constructed in Theorem 2.2. We know that
@@
We have an equivalence $\operatorname{MU}\otimes\mathbb{S}[x]/x^{e}\simeq f^{*}(\mathbb{S}[x]/x^{e})$ of $\mathbb{E}_{\infty}$-$\operatorname{MU}$-algebras.
@@
We have an equivalence of $\mathbb{E}_{3}$-$\operatorname{MU}$-algebras $A\simeq f_{*}(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}^{\mathbb{E}_{3}}\operatorname{MU}[x]/x^{e})$.
@@
It suffice to prove that $g_{*}g^{*}(\operatorname{BP}\langle n\rangle)\simeq\operatorname{BP}\langle n\rangle\otimes^{\mathbb{E}_{3}}_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$, which is clear an equivalence of $\operatorname{MU}$-modules. We only need to produce $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra map $\operatorname{BP}\langle n\rangle\otimes^{\mathbb{E}_{3}}_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}\to g_{*}g^{*}(\operatorname{BP}\langle n\rangle)$. According to [Lur17, Proposition 7.1.2.6], [Lur09, Proposition 5.2.2.8], we have a natural transformation of lax symmetrical monoidal functor $\operatorname{id}\to g_{*}g^{*}$, hence we have an adjunction pair between category of algebras
@@
One can also use the strategy in [HW22, §2] to construct an $\operatorname{MU}[x]$-algebra form of $\operatorname{BP}\langle n\rangle[x]/x^{e}$, i.e. by computing $\mathscr{U}^{(m)}_{\operatorname{MU}[x]}(\operatorname{BP}\langle n\rangle[x]/x^{e})$. It seems that one could only get an $\mathbb{E}_{1}$-form, due to computational difficulties. And we have following commutative diagram.
@@
For any spectrum $A,$ We define $\Lambda_{A}(x)$ to be the spectrum $A\otimes(\mathbb{S}\otimes_{\mathbb{S}[x]}\mathbb{S})$ where $\mathbb{S}$ view as an $\mathbb{S}[x]$-module via Construction 2.3 when $e=1$. If moreover, $A\in\operatorname{Alg}_{\mathbb{E}_{n}},$ where $1\leq n\leq\infty$. By the same argument, this gives rise an $\mathbb{E}_{n}$-algebra $\Lambda_{A}(x)$ in $\operatorname{Sp}.$ For instance, $\operatorname{MU}\in\operatorname{CAlg},$ then $\Lambda_{\operatorname{MU}}(x)$ is also a commutative algebra. Using Tor spectral sequence, one computes the homotopy group of $\Lambda_{\mathbb{S}}(x)$ as $\Lambda_{\pi_{*}\mathbb{S}}(x),$ where $|x|=1,$ see for instance [Ang08, Proposition 3.6]. We learn that $\pi_{*}(\Lambda_{\mathbb{S}}(x))$ is a free $\pi_{*}(\mathbb{S})$-module. Then if $A\in\operatorname{Alg}_{\mathbb{E}_{2}}$, we have isomorphism $\pi_{*}\Lambda_{A}(x)\simeq\Lambda_{\pi_{*}A}(x)$.
@@
where the $y$ lies in degree $2j\in 2\mathbb{N}_{>0}$. The commutative algebra $\operatorname{MU}[y]$ is very different from the Construction 2.3 unless $j=0$. We construct the $\mathbb{E}_{3}$-algebra form of $\operatorname{BP}\langle n\rangle$ by induction on $n.$ When $n=-1,0$, the $\mathbb{E}_{3}$-form (actually $\mathbb{E}_{\infty}$) of $\operatorname{BP}\langle n\rangle$ can be constructed via Postnikov truncation and note that [Lur17, Theorem 7.1.3.15]. Suppose we have an $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra form of $\operatorname{BP}\langle n\rangle$, we will inductively construct an $\mathbb{E}_{3}$-$\operatorname{MU}[y]/y^{k}$-algebra form $B_{k}$ using the square zero extension $\operatorname{MU}[y]/y^{k}\to\operatorname{MU}[y]/y^{k-1}$ and some information of the cotangent complex, where $|y|=2(p^{n+1}-1)$. More precisely, we have the following fact.
@@
So far, we have constructed several types of $\mathbb{E}_{3}$-algebras. Basically, all of them can be viewed as truncated polynomial algebras over $\operatorname{BP}\langle n\rangle$ for some $n$, but the polynomial generators may not live in the degree 00. The form in Proposition 2.7, $\operatorname{BP}\langle n\rangle[x]/x^{e}$ is just a finite direct sum of $\operatorname{BP}\langle n\rangle$ (note that $\operatorname{BP}\langle n\rangle[x]/x^{e}\simeq\operatorname{BP}\langle n\rangle_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$ as spectra). The form in Proposition 2.16 is equivalent to $\operatorname{BP}\langle n\rangle\bigoplus\Sigma^{1}\operatorname{BP}\langle n\rangle$ as $\operatorname{BP}\langle n\rangle$-module. The form in Proposition 2.22 is finite as $\operatorname{BP}\langle n\rangle$-module. Because by construction, the cofiber $\operatorname{cofib}(B_{k}\to B_{k-1})\simeq\Sigma^{k(2p^{n+1}-2)}\operatorname{BP}\langle n\rangle$, thus $\operatorname{BP}\langle n+1\rangle/v_{n+1}^{e}$ is a finite $\operatorname{BP}\langle n-1\rangle$-module. Since $\operatorname{MU}[y]/y^{e}$ is finite as $\operatorname{MU}$-module, we therefore conclude that the forms in Proposition 2.25 are finite as $\operatorname{BP}\langle n\rangle$-modules. We thus have proved following fact.
@@
In the following, we will use the conventions in Convention 2.12, 2.19, we will use the name of the form constructed in the Proposition 2.22,2.25. For instance, we will refer to $\operatorname{BP}\langle n\rangle/v_{n}^{e}$ as both algebra and $\operatorname{MU}$-algebra in Proposition 2.22. We now present some computational results of these algebras, before that, we present a well known lemma.
@@
There is an spectral sequence
@@
Let $A_{*}=\operatorname{BP}\langle n\rangle_{*}[x]/x^{e}$. We can constructed a ring map
@@
by choosing lifts of classes $\gamma_{p^{i}}(\sigma^{2}v_{n+1}),\gamma_{p^{i}}(\sigma^{2}x_{j}),\gamma_{p^{i}}(\sigma^{2}x^{e})$ in the Tor spectral sequence. Since the $\operatorname{BP}\langle n\rangle[x]/x^{e}\to\operatorname{BP}\langle 0\rangle[x]/x^{e}$ has the fiber lies in $\operatorname{Mod}_{\operatorname{MU}[x]}^{\geq|v_{1}|}$, we therefore conclude that the fiber of $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e}/\operatorname{MU}[x])\to\operatorname{THH}(\mathbb{Z}_{(p)}[x]/x^{e}/\operatorname{MU}[x])$ lies in $\operatorname{Mod}_{\operatorname{MU}[x]}^{\geq|v_{1}|}$. Assume $p<|v_{1}|=2p-2$, then $\operatorname{THH}_{p}(\mathbb{Z}_{(p)}[x]/x^{e}/\operatorname{MU}[x])\simeq\operatorname{THH}_{p}(\operatorname{BP}\langle n\rangle[x]/x^{e}/\operatorname{MU}[x])$, then according to the functoriality of Tor spectral sequence, we conclude that $p\gamma_{1}^{p}(\sigma^{2}x^{e})=\gamma_{p}(\sigma^{2}x^{e})$.

### v3

@@ 1.5. Notations and conventions
We use the language of $\infty$-categories developed in [LUR09], and the theory of higher algebra developed in [LUR17], for convenience. We denote the category of spectra by $\operatorname{Sp}$, and category of spaces by $\mathscr{S}$. And for an $\mathbb{E}_{n}$-monoidal category $\mathcal{C}$, $n\geq 2$, the category of $\mathbb{E}_{m}$-algebras over $\mathcal{C}$ is denoted by $\operatorname{Alg}_{\mathbb{E}_{m}}(\mathcal{C})$, where $0\leq m\leq n$. For $R\in\operatorname{Alg}_{\mathbb{E}_{n}}(\operatorname{Sp})$, we use $\operatorname{Alg}^{(m)}_{R}$ to denote $\operatorname{Alg}_{\mathbb{E}_{m}}(\operatorname{Mod}_{R})$. When $m=1,$ we write $\operatorname{Alg}_{R}^{(1)}$ as $\operatorname{Alg}_{R}$, when $m=\infty$, we write $\operatorname{Alg}_{R}^{(\infty)}$ as $\operatorname{CAlg}_{R}$. We use the symbol $\otimes^{\mathbb{E}_{m}}_{R}$ to indicate the tensor product in $\operatorname{Alg}^{(m)}_{R}$. When $R$ is the unit $\mathbb{S}\in\operatorname{Sp}$, we simply omit $R$. We refer to an object in $\operatorname{Alg}^{(m)}_{R}$ as an $\mathbb{E}_{m}$-$R$-algebra. When $m=1$, we simply refer to an object in $\operatorname{Alg}_{R}$ as an $R$-algebra. When $m=\infty$, we refer to an object in $\operatorname{CAlg}_{R}$ as a commutative $R$-algebra. When $R=\mathbb{S}$, we simply omit the word ’$R$’, or we call it $\mathbb{E}_{m}$-ring. A (discrete) ring simply indicate a classical ring. Let $R$ be a ring, we use $\otimes^{\mathbb{L}}$ to denote the derived tensor product. We use $\Lambda_{R}(x_{1},\ldots,x_{n})$ to denote the exterior algebra over $R$ generated by $x_{1},\ldots,x_{n}$. And we use $\Gamma_{R}(x_{1},\ldots,x_{n})$ to denote the divided power algebra generated by $x_{1},\ldots,x_{n}$, which sometimes we also write as $R\langle x_{1},\ldots,x_{n}\rangle$. For a stable $\infty$-category $\mathcal{C}$ with a $t$-structure, we use $\tau_{\leq n},\tau_{\geq n}$ to denote the truncation functors. When $\mathcal{C}=\operatorname{Sp}$, $X\in\operatorname{Sp}_{\leq n}$, if the localization $X\to\tau_{\leq n}X$ is an equivalence, and $X\in\operatorname{Sp}_{\geq m}$, if $\tau_{\geq m}X\to X$ is an equivalence. When $X\in\operatorname{Sp}_{\leq n}$, for some $n$, we say $X$ is bounded above, and $X\in\operatorname{Sp}_{\geq m}$, for some $m$, we say that $X$ is bounded below. If there are $a,b$ such that $X\in\operatorname{Sp}_{[a,b]}=\operatorname{Sp}_{\geq a}\cap\operatorname{Sp}_{\leq b}$, then $X$ is said to be bounded. Let $\mathcal{C}$ be a symmetrical monoidal category, and $A\in\operatorname{Alg}(\mathcal{C})$. We use [NS18, Definition III.2.3], to define $\operatorname{THH}(A)$, for instance, $A\in\operatorname{Alg}$, $\operatorname{THH}(A)=A\otimes_{A\otimes A^{\textup{op}}}A$.
@@ 1.6. Acknowledgments
The author would like to thank his advisor Guozhen Wang for carefully reading the early drafts. The author would like to thank Chris Brav for helping to improve the writing of this article. The author would like especially to thank Lars Hesselholt, for suggesting the author Proposition 4.9, and telling the author Proposition 4.19. The author would like to thank Gabriel Knoll for detailed feedback. The author would like to thank Oliver Braunling and John Rognes for comments. Finally, the author thanks Jingbang Guo and Wei Yang, for useful conversations. The author is partially supported by Grant NSFC-12226002. The author sincerely thanks the IAS/PCMI Summer School and IWoAT for their inspiring programs. The author benefited greatly from these two wonderful conferences, during which part of this work was completed.
@@ 2. Truncated polynomials and infinitesimal thickening
We define $\mathbb{S}[x]:=\Sigma^{\infty}_{+}\mathbb{N}.$ In a same fashion, we may consider the truncated version. For $e\in\mathbb{N}_{\geq 1}$, $M_{e}:=\{1,x,\ldots,x^{e-1}\}\simeq\mathbb{N}/([e,\infty)\cap\mathbb{N})$ where the multiplication is defined in an obvious way. The multiplication of $M_{e},\mathbb{N}$ are strictly commutative. Unwinding the definitions in [LUR17, Definition 2.4.2.1, Remark 2.4.2.2], we learn that $\mathbb{N},M_{e}$ are commutative monoid objects in $\mathscr{S}$, and hence commutative algebra objects in $\mathscr{S}$. Let $\mathbb{S}[x]/x^{e}:=\Sigma^{\infty}_{+}M_{e}\in\operatorname{Sp}.$ There are $\mathbb{E}_{\infty}$-ring maps
@@
$\mathbb{N}\to\mathbb{N}/([e,\infty)\cap\mathbb{N})$
@@
$\mathbb{N}/[e,\infty)\to\mathbb{N}/([1,\infty)\cap\mathbb{N})$
@@
As a spectrum $\mathbb{S}[x]$ is nothing but a direct sum of $\mathbb{S}$, and this is also true for $\mathbb{S}[x]/x^{e},$ using the multiplication on $\mathbb{N},M_{e},$ we can get
@@
In general, let $M$ be a commutative monoid, one can define $A[M]:=A\otimes\Sigma^{\infty}_{+}M$, for any spectrum $A.$ For example, $M$ is a sub monoid of $\mathbb{N}^{\times m}$, or a quotient monoid of $\mathbb{N}^{\times m}.$ In a upcoming paper, we will consider $\operatorname{BP}\langle n\rangle[x,y]/(xy)$.
@@
There is an $\mathbb{E}_{3}$-algebra $A$ such that $A$ is the image of an $\mathbb{E}_{3}$-$\operatorname{MU}[x]$-algebra form of $\operatorname{BP}\langle n\rangle[x]/x^{e}$ under the forgetful functor.
@@
We let $A$ denote the $\mathbb{E}_{3}$-algebra in Proposition 2.7 and let $B$ denote the underlying $\mathbb{E}_{3}$-algebra of the $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra $\operatorname{BP}\langle n\rangle$ constructed in Theorem 2.2. We know that
@@
We have an equivalence of $\mathbb{E}_{\infty}$-$\operatorname{MU}$-algebras $\operatorname{MU}\otimes\mathbb{S}[x]/x^{e}\simeq f^{*}(\mathbb{S}[x]/x^{e})$.
@@
We have an equivalence of $\mathbb{E}_{3}$-rings $A\simeq f_{*}(\operatorname{BP}\langle n\rangle\otimes_{\operatorname{MU}}^{\mathbb{E}_{3}}\operatorname{MU}[x]/x^{e})$.
@@
Because $A\simeq f_{*}g_{*}(g^{*}(\operatorname{BP}\langle n\rangle))$. It suffice to prove that $g_{*}g^{*}(\operatorname{BP}\langle n\rangle)\simeq\operatorname{BP}\langle n\rangle\otimes^{\mathbb{E}_{3}}_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$ is an equivalence of $\mathbb{E}_{3}$-$\operatorname{MU}$-algebras, which is clearly an equivalence of $\operatorname{MU}$-modules. We only need to produce an $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra map $\operatorname{BP}\langle n\rangle\otimes^{\mathbb{E}_{3}}_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}\to g_{*}g^{*}(\operatorname{BP}\langle n\rangle)$. According to [LUR17, Proposition 7.1.2.6], [LUR09, Proposition 5.2.2.8], we have a natural transformation of lax symmetrical monoidal functors $\operatorname{id}\to g_{*}g^{*}$, hence we have an adjunction pair
@@
One can also use the strategy in [HW22, §2] to construct an $\operatorname{MU}[x]$-algebra form of $\operatorname{BP}\langle n\rangle[x]/x^{e}$, i.e. by computing $\mathscr{U}^{(m)}_{\operatorname{MU}[x]}(\operatorname{BP}\langle n\rangle[x]/x^{e})$. It seems that one could only get an $\mathbb{E}_{1}$-form, due to computational difficulties. We have following commutative diagram.
@@
For any spectrum $A,$ We define $\Lambda_{A}(x)$ to be the spectrum $A\otimes(\mathbb{S}\otimes_{\mathbb{S}[x]}\mathbb{S})$ where $\mathbb{S}$ view as an $\mathbb{S}[x]$-module via Construction 2.3 when $e=1$. If moreover, $A\in\operatorname{Alg}_{\mathbb{E}_{n}},$ where $1\leq n\leq\infty$. By the same argument, this gives rise an $\mathbb{E}_{n}$-algebra $\Lambda_{A}(x)$ in $\operatorname{Sp}.$ For instance, $\operatorname{MU}\in\operatorname{CAlg},$ then $\Lambda_{\operatorname{MU}}(x)$ is also a commutative algebra. Using the Tor spectral sequence, one computes the homotopy group of $\Lambda_{\mathbb{S}}(x)$ as $\Lambda_{\pi_{*}\mathbb{S}}(x),$ where $|x|=1,$ see for instance [ANG08, Proposition 3.6]. We learn that $\pi_{*}(\Lambda_{\mathbb{S}}(x))$ is a free $\pi_{*}(\mathbb{S})$-module. Then if $A\in\operatorname{Alg}_{\mathbb{E}_{2}}$, we have isomorphism $\pi_{*}\Lambda_{A}(x)\simeq\Lambda_{\pi_{*}A}(x)$.
@@
where the $y$ lies in degree $2j\in 2\mathbb{N}_{>0}$. The commutative algebra $\operatorname{MU}[y]$ is very different from the Construction 2.3 unless $j=0$. We construct the $\mathbb{E}_{3}$-algebra form of $\operatorname{BP}\langle n\rangle$ by induction on $n.$ When n=1,0n=-1,0, the $\mathbb{E}_{3}$-form (actually $\mathbb{E}_{\infty}$) of $\operatorname{BP}\langle n\rangle$ can be constructed via Postnikov truncation and note that [LUR17, Theorem 7.1.3.15]. Suppose we have an $\mathbb{E}_{3}$-$\operatorname{MU}$-algebra form of $\operatorname{BP}\langle n\rangle$, we will inductively construct an $\mathbb{E}_{3}$-$\operatorname{MU}[y]/y^{k}$-algebra form $B_{k}$ using the square zero extension $\operatorname{MU}[y]/y^{k}\to\operatorname{MU}[y]/y^{k-1}$ and some information of the cotangent complex, where $|y|=2(p^{n+1}-1)$. More precisely, we have the following fact.
@@
So far, we have constructed several types of $\mathbb{E}_{3}$-algebras. Basically, all of them can be viewed as truncated polynomial algebras over $\operatorname{BP}\langle n\rangle$ for some $n$, but the polynomial generators may not live in the degree 0. The form in Proposition 2.7, $\operatorname{BP}\langle n\rangle[x]/x^{e}$ is just a finite direct sum of $\operatorname{BP}\langle n\rangle$ (note that $\operatorname{BP}\langle n\rangle[x]/x^{e}\simeq\operatorname{BP}\langle n\rangle_{\operatorname{MU}}\operatorname{MU}[x]/x^{e}$ as spectra). The form in Proposition 2.16 is equivalent to $\operatorname{BP}\langle n\rangle\bigoplus\Sigma^{1}\operatorname{BP}\langle n\rangle$ as $\operatorname{BP}\langle n\rangle$-module. The form in Proposition 2.22 is finite as $\operatorname{BP}\langle n\rangle$-module. Because by construction, the cofiber $\operatorname{cofib}(B_{k}\to B_{k-1})\simeq\Sigma^{k(2p^{n+1}-2)}\operatorname{BP}\langle n\rangle$, thus $\operatorname{BP}\langle n+1\rangle/v_{n+1}^{e}$ is a finite $\operatorname{BP}\langle n\rangle$-module. Since $\operatorname{MU}[y]/y^{e}$ is finite as $\operatorname{MU}$-module, we therefore conclude that the forms in Proposition 2.25 are finite as $\operatorname{BP}\langle n\rangle$-modules. We thus have proved following fact.
@@
In the following, we will use the conventions in Convention 2.12, 2.19, we will use the name of the form constructed in the Proposition 2.22, 2.25. For instance, we will refer to $\operatorname{BP}\langle n\rangle/v_{n}^{e}$ as both an algebra and an $\operatorname{MU}$-algebra in Proposition 2.22. We now present some computational results of these algebras, before that, we present a well known lemma.
@@
There is a spectral sequence
@@
Let $A_{*}=\operatorname{BP}\langle n\rangle_{*}[x]/x^{e}$. We can construct a ring map
@@
by choosing lifts of classes $\gamma_{p^{i}}(\sigma^{2}v_{n+1}),\gamma_{p^{i}}(\sigma^{2}x_{j}),\gamma_{p^{i}}(\sigma^{2}x^{e})$ in the Tor spectral sequence. Since the fiber of $\operatorname{BP}\langle n\rangle[x]/x^{e}\to\operatorname{BP}\langle 0\rangle[x]/x^{e}$ lies in $\operatorname{Mod}_{\operatorname{MU}[x]}^{\geq|v_{1}|}$, we conclude that the fiber of $\operatorname{THH}(\operatorname{BP}\langle n\rangle[x]/x^{e}/\operatorname{MU}[x])\to\operatorname{THH}(\mathbb{Z}_{(p)}[x]/x^{e}/\operatorname{MU}[x])$ lies in $\operatorname{Mod}_{\operatorname{MU}[x]}^{\geq|v_{1}|}$. Assume $p<|v_{1}|=2p-2$, then $\operatorname{THH}_{p}(\mathbb{Z}_{(p)}[x]/x^{e}/\operatorname{MU}[x])\simeq\operatorname{THH}_{p}(\operatorname{BP}\langle n\rangle[x]/x^{e}/\operatorname{MU}[x])$, then according to the functoriality of Tor spectral sequence, we conclude that $p\gamma_{1}^{p}(\sigma^{2}x^{e})=\gamma_{p}(\sigma^{2}x^{e})$.